# The $s\pm$ pairing symmetry in the pressured La3Ni2O7 from electron-phonon coupling

### v1

@@ I Introduction
The search for a new family of high transition temperature (high-$T_{c}$) superconductors remains the central task in the condensed matter community, since the discovery of copper-based high-$T_{c}$ superconductorsBednorz and Müller (1986); Anderson (1987); Lee et al. (2006); Keimer et al. (2015). Drawing inspiration from the crystal and electronic structures of cuprates, the iron-based superconductorsKamihara et al. (2008); Daghofer et al. (2008); Raghu et al. (2008); Chubukov et al. (2008); Eschrig and Koepernik (2009); Stewart (2011) and the recent infinite layer nickelate superconductorsLi et al. (2019); Been et al. (2021); Lee and Pickett (2004); Hepting et al. (2020); Botana and Norman (2020); Wu et al. (2020); Jiang et al. (2020) are discovered. Notably, the discovery of superconducting property in bilayer Ruddlesden-Popper nickelate La3Ni2O7 (LNO)Sun et al. (2023), which exhibits a remarkable transition temperature $T_{c}\approx 80$ K under high pressure exceeding 14 GPa, has sparked intensive investigation into this novel superconducting materialHou et al. (2023); Zhang et al. (2024a); Wang et al. (2024); Luo et al. (2023); Lechermann et al. (2023); Gu et al. (2023); Yang et al. (2023a); Christiansson et al. (2023); Sakakibara et al. (2024); Shilenko and Leonov (2023); Shen et al. (2023); Liu et al. (2024); Wú et al. (2024); Chen et al. (2025); Cao and Yang (2024); Liu et al. (2023a); Zhang et al. (2023a); Lu et al. (2024); Zhang et al. (2024b); Oh and Zhang (2023); Liao et al. (2023); Qu et al. (2024); Yang et al. (2023b); Jiang et al. (2024); Zhang et al. (2023b); Huang et al. (2023); Tian et al. (2024); Qin and Yang (2023); Heier et al. (2024); Khasanov et al. (2024); Meng et al. (2024); Yang et al. (2024); Yi et al. (2024); Liu et al. ; Liu et al. (2023b); Geisler et al. (2024). Subsequent studiesHou et al. (2023); Zhang et al. (2024a); Wang et al. (2024) have further confirmed the observation of zero electrical resistance, solidifying LNO as a promising candidate for further exploration in the field of high-$T_{c}$ superconductivity. Increasing pressure causes the structure to undergo a transition from the orthorhombic structure of the Amam space group to a regular AA-stacking structure of the Fmmm space group. In contrast to the $d^{9}$ electronic configuration in cuprates, the normal valence band in LNO manifests the $d^{7.5}$ configurationSun et al. (2023); Zhang et al. (2024a). The density functional theory (DFT) further reveals that the band structure in the vicinity of the Fermi level under pressure is mainly dominated by the Ni-$d_{x^{2}-y^{2}}$ and Ni-$d_{3z^{2}-r^{2}}$ orbitalsSun et al. (2023). These different features imply the fundamental diverse pairing symmetry from the cuprates. Interestingly, in addition to superconductivity, the charge/spin density wave, and non-Fermi-liquid behavior are also suggested by transport measurementsSun et al. (2023); Liu et al. (2023b); Zhang et al. (2024a); Liu et al. .
@@
A fundamental question in the study of this new family of high-$T_{c}$ superconductors revolves around elucidating the underlying superconducting mechanism. Current theoretical efforts have predominantly focused on understanding superconductivity in terms of the unique band structure and electronic correlations present in this material. However, recent experimental findings suggest that electron-phonon coupling may play a pivotal role in determining the electronic properties. The ongoing debate among several research groups on the strength of electron-phonon couplingTalantsev and Chistyakov (2024); Li et al. (2024); You et al. (2025); Ouyang et al. (2024); Zhan et al. (2024) raises a critical question: What would be the pairing symmetry in pressurized LNO if the superconducting state is indeed mediated by electron-phonon ($e$-ph) coupling?
@@ II Model and method
The low-energy electronic structure of bilayer La3Ni2O7, as revealed by DFT calculations and experiments, is predominantly governed by the Ni-$d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$ orbitalsSun et al. (2023); Luo et al. (2023); Lechermann et al. (2023); Gu et al. (2023). Motivated by this orbital-selective character, we take a minimal bilayer two-orbital model adapted from Ref. Luo et al. (2023). This effective model successfully encapsulates the essential features of the low-energy band structure near the Fermi level, including orbital hybridization effects and interlayer coupling mechanisms. The model reads
@@
Here $T^{\eta=x^{2},z^{2}}$ represents the intralayer intraorbital hopping term, while $V_{\bf k}$ ($V^{\prime}_{\bf k}$) demonstrates the interorbital intralayer (interlayer) hopping. $t_{\perp}^{x^{2}}$ ($t_{\perp}^{z^{2}}$) denotes the interlayer hopping term in Ni-$d_{x^{2}-y^{2}}$ ($d_{3z^{2}-r^{2}}$) orbital. The detailed values of hopping parameters $t_{1}^{\eta}$, $t_{2}^{\eta}$, $t_{3}$, $t_{4}$ can be found in the referenceLuo et al. (2023). Fig. 1(a) presents the calculated band structures of the system, while panels (b)-(d) display the corresponding Fermi surfaces at three representative chemical potentials. In addition to the prominent high-$T_{c}$ superconducting regime near the critical doping level reported in Ref. Sun et al. (2023), we systematically investigate two characteristic doping scenarios: (i) a hole-doped regime with $\mu=-0.328eV$, where the Fermi surface topology exhibits a van Hove singularity, and (ii) an electron-doped regime with $\mu=0.1eV$. Fig. 1(b-d) reveals that there exist three Fermi surfaces: two hole-like pockets $\beta$ and $\gamma$ originating from the interlayer antibonding states of the Ni-$d_{x^{2}-y^{2}}$ and bonding state of $d_{3z^{2}-r^{2}}$ orbitals; one electron-like pocket formed by the interlayer bonding states of both $d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$ orbitals. Notably, the orbital-projection analysis demonstrates that the $\gamma$ pocket predominantly derives its spectral weight from the $d_{3z^{2}-r^{2}}$ orbital character.
Figure 1: The orbital-weight band structureLuo et al. (2023) (a) and Fermi surfaces (b-d) of the bilayer tight-binding model in the pressured LNO. (b) labels the Fermi surface with $\mu=-0.328eV$ where it manifest a van Hove singularity. (c) and (d) illustrates the Fermi surface with $\mu=0eV$ and $0.1eV$, respectively.
@@
Figure 2: The pairing gap in the full-coupling case with different parameters. (a), (b) and (c) illustrates the numerical result with only interlayer coupling $g_{2}^{f}=0.12$ for $\mu=-0.328eV$, $0eV$ and $0.1eV$ respectively. (d), (e) and (f) illustrates the numerical result with only intralayer coupling $g_{1}^{f}=0.06$ for $\mu=-0.328eV$, $0eV$ and $0.1eV$ respectively.
@@
×$\times c_{\beta^{\prime}s^{\prime}}(-{\bf k}^{\prime})c_{\beta^{\prime}s}({\bf k}^{\prime})$ (8)
@@
near the superconducting transition temperature $T_{c}$. The superconducting transition temperature $T_{c}$ is the highest temperature such that the matrix $\chi_{{\bf k},{\bf k}^{\prime}}^{(\beta,\beta^{\prime})}$ yields eigenvalue 1 in the discrete momentum and band spanned spaceChou et al. (2021).
Figure 4: The pairing gap with fixed interlayer coupling $g_{2}^{f}=0.12$ while varying intralayer coupling $g_{1}^{f}=0.01$ (a) and $0.07$ (b) in the full-coupling case with chemical potential $\mu=0$.
@@ III Superconducting state for the full-coupling case
×$\times U_{\eta\alpha\beta^{\prime}}({\bf k}^{\prime}),$ (14)
@@
×$\times U_{t\alpha\beta^{\prime}}({\bf k}^{\prime}).$ (15)
@@
To clearly understand the role of intralayer and interlayer coupling in the pairing symmetry, we separately examine the intralayer and interlayer couplings and plot the resulting pairing symmetries for various doping levels in Fig. 2. Fig. 2(a-c) show that the pairing symmetry induced by the interlayer BCS coupling in the full-coupling case results in pairing gaps with the same sign in the $\alpha$ and $\gamma$ pockets, while the gap between the $\alpha$ and $\beta$ pockets has the opposite sign, for fixed parameters $g_{1}^{f}=0$ and $g_{2}^{f}=0.12$ under various doping conditions. Thus, our numerical calculations for the full-coupling case with only the interlayer coupling indicate that the superconducting state hosts an $s\pm$-wave pairing symmetry.
Figure 5: The pairing gap in the half-coupling case with different parameters. (a), (b) and (c) illustrates the numerical result with only interlayer coupling $g_{2}^{f}=0.18$ for $\mu=-0.328eV$, $0eV$ and $0.1eV$ respectively. (d) The result with parameters $g_{1}^{f}=0.11$, $g_{2}^{f}=0$ and $\mu=-0.328eV$. (e) The result with parameters $g_{1}^{f}=0.17$, $g_{2}^{f}=0$ and $\mu=0eV$. (f) The result with parameters $g_{1}^{f}=0.195$, $g_{2}^{f}=0$ and $\mu=0.1eV$.
Furthermore, to address the origin of the $s\pm$ pairing symmetry induced by interlayer coupling, we plot the pairing interactions in the band space in Fig. 3(a-c). As illustrate in Fig. 3(a,b), the pairing interaction $g^{\alpha,\alpha}$ and $g^{\alpha\gamma}$ indicate negative Josephson coupling both within $\alpha$ intra-Fermi pocket and between $\alpha$ and $\gamma$ Fermi pockets, intriguing to pairing gaps with the same sign within the same Fermi pockets and between $\alpha$ and $\gamma$ Fermi pockets. Meanwhile, the negative $g^{\alpha,\beta}$ will favor the opposite sign pairing gap between $\alpha$ and $\gamma$ Fermi pockets. Conversely, the pure intralayer coupling scenario will sustains the conventional $s++$-wave superconducting state with uniform pairing gap across all pockets as shown in Fig. 2(d-e) with fixed parameter $g_{1}^{f}=0.06$ and $g_{2}^{f}=0$.
@@
Figure 6: The pairing gap in the half-coupling case with fixed interlayer coupling $g^{h}_{2}=0.18$ while varying intralayer coupling $g_{1}^{h}=0.165$ (a) and $0.18$ (b) with chemical potential $\mu=0$
@@ IV Pairing symmetry in half-coupling case
×$\times U_{\eta x^{2}\beta^{\prime}}({\bf k}^{\prime}),$ (16)
@@
Following the methodology employed in the full-coupling scenario, we systematically investigate the individual effects of intralayer and interlayer couplings. By numerically solving the self-consistent equations Eq. (9) and (13) with the interaction $g^{\beta,\beta^{\prime}}$, we plot the pairing gaps in Fig. 5. As shown in Fig. 5(a-c) with fixed parameters $g_{1}^{h}=0$ and $g_{2}^{h}=0.18$ with several doping conditions, the interlayer coupling also favors the $s\pm$-wave pairing symmetry as the same as in the full-coupling case. In Fig. 5(d-e), we use the different intralayer coupling strengths $g_{2}^{h}$ to drive the superconducting state and obtain the pairing gap under different doping conditions. Then we reach the same conclusion in the half-coupling case as in the full-coupling scenario, the intralayer coupling $g_{2}^{h}$ also tends to induce the conventional $s++$-wave pairing symmetry, due to the same reasons as discussed in the full-coupling case.
Figure 7: The pairing gaps with fixed $\mu=0eV$. (a) and (b) marks the pairing gap in the full-coupling case with fixed parameters $g_{1}^{f}=0.07$ and $g_{2}^{f}=0.12$ while varying pair hoping $g_{P}=-0.02$ and $0.02$, respectively. (c) and (d) illustrates the pairing gap in the half-coupling case with fixed parameters $g_{1}^{h}=0.18$ and $g_{2}^{h}=0.18$ while varying pair hoping $g_{P}=-0.02$ and $0.02$, respectively.
@@ V The role of pair hopping interaction
×$\times c_{\beta^{\prime}s^{\prime}}(-{\bf k}^{\prime})c_{\beta^{\prime}s}({\bf k}^{\prime})$ (19)
@@
×$\times U_{\eta z^{2}\beta^{\prime}}({\bf k}^{\prime}).$ (20)
@@
Compared to Fig.4(b), the numerical results presented in Fig.7(a) in the full-coupling model indicate that a negative pair-hopping interaction $g_{P}$ tends to drive the superconducting state toward an alternative $s\pm$-wave state. In this state, the pairing gap on the $\alpha$ Fermi surface shares the same sign as the gap on the $\beta$ Fermi surface but has the opposite sign compared to the gap on the $\gamma$ Fermi surface. In contrast, a positive $g_{P}$ in the full-coupling case favors the $s\pm$-wave pairing symmetry. The preference for $s++$-wave symmetry by negative and $s\pm$-wave symmetry by positive $g_{P}$ can be attributed to the fact that the Fermi surface is primarily composed of $d_{3z^{2}-r^{2}}$ orbital weight, while both the $\alpha$ and $\beta$ Fermi surfaces have significant contributions from the $d_{x^{2}-y^{2}}$ orbital. Fig. 7 (c) and (d) lead to the conclusion that in the half-coupling case the pair-hopping strength $g_{P}$ plays a similar role as in the full-coupling case.

### v2

# The $s\pm$ pairing symmetry from electron-phonon coupling in La3Ni2O7 under pressure

@@ I Introduction
The search for a new family of high transition temperature (high-$T_{c}$) superconductors remains the central task in the condensed matter community, since the discovery of copper-based high-$T_{c}$ superconductors[1, 2, 3, 4]. Drawing inspiration from the crystal and electronic structures of cuprates, the iron-based superconductors[5, 6, 7, 8, 9, 10] and the recent infinite layer nickelate superconductors[11, 12, 13, 14, 15, 16, 17] are discovered. Notably, the discovery of superconducting property in bilayer Ruddlesden-Popper nickelate La3Ni2O7 (LNO)[18], which exhibits a remarkable transition temperature $T_{c}\approx 80$ K under high pressure exceeding 14 GPa, has sparked intensive investigation into this novel superconducting material[19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Subsequent studies[19, 20, 21] have further confirmed the observation of zero electrical resistance, solidifying LNO as a promising candidate for further exploration in the field of high-$T_{c}$ superconductivity. Increasing pressure causes the structure to undergo a transition from the orthorhombic structure of the Amam space group to a regular AA-stacking structure of the Fmmm space group. In contrast to the $d^{9}$ electronic configuration in cuprates, the normal valence band in LNO manifests the $d^{7.5}$ configuration[18, 20]. The density functional theory (DFT) further reveals that the band structure in the vicinity of the Fermi level under pressure is mainly dominated by the Ni-$d_{x^{2}-y^{2}}$ and Ni-$d_{3z^{2}-r^{2}}$ orbitals[18]. These different features imply the fundamental diverse pairing symmetry from the cuprates. Interestingly, in addition to superconductivity, the charge/spin density wave, and non-Fermi-liquid behavior are also suggested by transport measurements[18, 53, 20, 30].
@@
A fundamental question in the study of this new family of high-$T_{c}$ superconductors revolves around elucidating the underlying superconducting mechanism. Current theoretical efforts have predominantly focused on understanding superconductivity in terms of the unique band structure and electronic correlations present in this material. However, recent experimental findings suggest that electron-phonon coupling may play a pivotal role in determining the electronic properties. The ongoing debate among several research groups on the strength of electron-phonon coupling[60, 61, 62, 63, 64] (EPC) raises a critical question: What would be the pairing symmetry in pressurized LNO if the superconducting state is indeed mediated by electron-phonon ($e$-ph) coupling?
@@ II Model and method
The low-energy electronic structure of bilayer La3Ni2O7, as revealed by DFT calculations and experiments, is predominantly governed by the Ni-$d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$ orbitals[18, 22, 23, 24]. Motivated by this orbital-selective character, we take a minimal bilayer two-orbital model adapted from Ref. [22]. This effective model successfully encapsulates the essential features of the low-energy band structure near the Fermi level, including orbital hybridization effects and interlayer coupling mechanisms. The model reads
@@
Here $T^{\eta=x^{2},z^{2}}$ represents the intralayer intraorbital hopping term, while $V_{\bf k}$ ($V^{\prime}_{\bf k}$) demonstrates the interorbital intralayer (interlayer) hopping. $t_{\perp}^{x^{2}}$ ($t_{\perp}^{z^{2}}$) denotes the interlayer hopping term in Ni-$d_{x^{2}-y^{2}}$ ($d_{3z^{2}-r^{2}}$) orbital. The detailed values of hopping parameters $t_{1}^{\eta}$, $t_{2}^{\eta}$, $t_{3}$, $t_{4}$ can be found in the reference[22]. Fig. 1(a) presents the calculated band structures of the system, while panels (b)-(d) display the corresponding Fermi surfaces at three representative chemical potentials. In addition to the prominent high-$T_{c}$ superconducting regime near the critical doping level reported in Ref. [18], we systematically investigate two characteristic doping scenarios: (i) a hole-doped regime with $\mu=-0.328eV$, where the Fermi surface topology exhibits a van Hove singularity, and (ii) an electron-doped regime with $\mu=0.1eV$. Fig. 1(b-d) reveals that there exist three Fermi surfaces: two hole-like pockets $\beta$ and $\gamma$ originating from the interlayer antibonding states of the Ni-$d_{x^{2}-y^{2}}$ and bonding state of $d_{3z^{2}-r^{2}}$ orbitals; one electron-like pocket formed by the interlayer bonding states of both $d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$ orbitals. Notably, the orbital-projection analysis demonstrates that the $\gamma$ pocket predominantly derives its spectral weight from the $d_{3z^{2}-r^{2}}$ orbital character.
Figure 1: The orbital-weight band structure[22] (a) and Fermi surfaces (b-d) of the bilayer tight-binding model in the pressured LNO. (b) labels the Fermi surface with $\mu=-0.328eV$ where it manifest a van Hove singularity. (c) and (d) illustrates the Fermi surface with $\mu=0eV$ and $0.1eV$, respectively.
@@
Figure 2: The pairing gap in the full-coupling case with different parameters. (a), (b) and (c) illustrates the numerical result with only intralayer coupling $g_{2}^{f}=0.12$ for $\mu=-0.328eV$, $0eV$ and $0.1eV$ respectively. (d), (e) and (f) illustrates the numerical result with only interlayer coupling $g_{1}^{f}=0.12$ for $\mu=-0.328eV$, $0eV$ and $0.1eV$ respectively.
@@
$\times c_{\beta^{\prime}s^{\prime}}(-{\bf k}^{\prime})c_{\beta^{\prime}s}({\bf k}^{\prime})$ (8)
@@
$\Delta^{\beta}({\bf k})=\frac{1}{A}\sum_{\beta^{\prime},{\bf k}^{\prime}}\chi^{\beta,\beta^{\prime}}({\bf k},{\bf k}^{\prime})\Delta^{\beta^{\prime}}({\bf k}^{\prime}),$ (13)
with
@@
near the superconducting transition temperature $T_{c}$. The superconducting transition temperature $T_{c}$ is the highest temperature such that the matrix $\chi_{{\bf k},{\bf k}^{\prime}}^{(\beta,\beta^{\prime})}$ yields eigenvalue 1 in the discrete momentum and band spanned space[65].
Figure 4: The pairing gap with fixed interlayer coupling $g_{2}^{f}=0.12$ while varying intralayer coupling $g_{1}^{f}=0.006$ (a) and $0.072$ (b) in the full-coupling case with chemical potential $\mu=0$.
@@ III Superconducting state for the full-coupling case
$\times U_{\eta\alpha\beta^{\prime}}({\bf k}^{\prime}),$ (15)
@@
$\times U_{t\alpha\beta^{\prime}}({\bf k}^{\prime}).$ (16)
@@
To clearly understand the role of intralayer and interlayer coupling in the pairing symmetry, we separately examine the intralayer and interlayer couplings and plot the resulting pairing symmetries for various doping levels in Fig. 2. Fig. 2(d-f) show that the pairing symmetry induced by the interlayer BCS coupling in the full-coupling case results in pairing gaps with the same sign in the $\alpha$ and $\gamma$ pockets, while the gap between the $\alpha$ and $\beta$ pockets has the opposite sign, for fixed parameters $g_{1}^{f}=0$ and $g_{2}^{f}=0.12$ under various doping conditions. Thus, our numerical calculations for the full-coupling case with only the interlayer coupling indicate that the superconducting state hosts an $s\pm$-wave pairing symmetry.
Figure 5: The pairing gap in the half-coupling case with different parameters. (a), (b) and (c) illustrates the numerical result with only intralayer coupling $g_{1}^{f}=0.18$ for $\mu=-0.328eV$, $0eV$ and $0.1eV$ respectively. (d), (e) and (f) illustrates the numerical result with only interlayer coupling $g_{2}^{f}=0.18$ for $\mu=-0.328eV$, $0eV$ and $0.1eV$ respectively.
Furthermore, to address the origin of the $s\pm$ pairing symmetry induced by interlayer coupling, we plot the pairing interactions in the band space in Fig. 3(a-c). As illustrate in Fig. 3(a,b), the pairing interaction $g^{\alpha,\alpha}$ and $g^{\alpha\gamma}$ indicate negative Josephson coupling both within $\alpha$ intra-Fermi pocket and between $\alpha$ and $\gamma$ Fermi pockets, intriguing to pairing gaps with the same sign within the same Fermi pockets and between $\alpha$ and $\gamma$ Fermi pockets. Meanwhile, the negative $g^{\alpha,\beta}$ will favor the opposite sign pairing gap between $\alpha$ and $\gamma$ Fermi pockets. Conversely, the pure intralayer coupling scenario will sustains the conventional $s++$-wave superconducting state with uniform pairing gap across all pockets as shown in Fig. 2(a-c) with fixed parameter $g_{1}^{f}=0.12$ and $g_{2}^{f}=0$.
@@
Figure 6: The pairing gap in the half-coupling case with fixed interlayer coupling $g^{h}_{2}=0.18$ while varying intralayer coupling $g_{1}^{h}=0.016$ (a) and $0.05$ (b) with chemical potential $\mu=0$
@@ IV Pairing symmetry in half-coupling case
$\times U_{\eta x^{2}\beta^{\prime}}({\bf k}^{\prime}),$ (17)
@@
Following the methodology employed in the full-coupling scenario, we systematically investigate the individual effects of intralayer and interlayer couplings. By numerically solving the self-consistent equations Eq. (9) and (14) with the interaction $g^{\beta,\beta^{\prime}}$, we plot the pairing gaps in Fig. 5. As shown in Fig. 5(d-f) with fixed parameters $g_{1}^{h}=0$ and $g_{2}^{h}=0.18$ with several doping conditions, the interlayer coupling also favors the $s\pm$-wave pairing symmetry as the same as in the full-coupling case. In Fig. 5(a-c), we use the different intralayer coupling strengths $g_{2}^{h}$ to drive the superconducting state and obtain the pairing gap under different doping conditions. Then we reach the same conclusion in the half-coupling case as in the full-coupling scenario, the intralayer coupling $g_{2}^{h}$ also tends to induce the conventional $s++$-wave pairing symmetry, due to the same reasons as discussed in the full-coupling case.
Figure 7: The pairing gaps with fixed $\mu=0eV$. (a) and (b) marks the pairing gap in the full-coupling case with fixed parameters $g_{1}^{f}=0.006$ and $g_{2}^{f}=0.12$ while varying pair hoping $g_{P}=-0.1$ and $0.1$, respectively. (c) and (d) illustrates the pairing gap in the half-coupling case with fixed parameters $g_{1}^{h}=0.016$ and $g_{2}^{h}=0.18$ while varying pair hoping $g_{P}=-0.1$ and $0.1$, respectively.
@@ V The role of pair hopping interaction
$\times c_{\beta^{\prime}s^{\prime}}(-{\bf k}^{\prime})c_{\beta^{\prime}s}({\bf k}^{\prime})$ (20)
@@
$\times U_{\eta z^{2}\beta^{\prime}}({\bf k}^{\prime}).$ (21)
@@
Compared to Fig.4(a), the numerical results presented in Fig.7(a) in the full-coupling model indicate that a negative pair-hopping interaction $g_{P}$ tends to drive the superconducting state toward an alternative $s\pm$-wave state. In this state, the pairing gap on the $\alpha$ Fermi surface shares the same sign as the gap on the $\beta$ Fermi surface but has the opposite sign compared to the gap on the $\gamma$ Fermi surface. In contrast, a positive $g_{P}$ in the full-coupling case favors the conventional $s++$-wave pairing symmetry. The preference for $s\pm$-wave symmetry by negative and $s++$-wave symmetry by positive $g_{P}$ can be attributed to the fact that the Fermi surface is primarily composed of $d_{3z^{2}-r^{2}}$ orbital weight, while both the $\alpha$ and $\beta$ Fermi surfaces have significant contributions from the $d_{x^{2}-y^{2}}$ orbital. Fig. 7 (c) and (d) lead to the conclusion that in the half-coupling case the pair-hopping strength $g_{P}$ plays a similar role as in the full-coupling case.